# A Survey on Self-Evolution of Large Language Models

### v1

@@ 6.2 In-Context
This approach utilizes an external module to collect, update, and retrieve past experiences and knowledge, enabling models to access a rich pool of insights and achieve better results without updating model parameters. The external memory mechanism is common in AI Agent systems Xu et al. (2023b); Qian et al. (2024); Wang et al. (2024d). This section provides a detailed overview of the latest methods for updating external memory, emphasizing the aspects of memory Content and Updating Operations, and summarized in Table 2.
@@
Content: External memory mainly stores two type of content: past experiences and reflected rationale, each serving distinct purposes. For instance, past experience provide valuable historical context, serving as a guiding force toward achieving improved outcomes. MoT Li and Qiu (2023) archives filtered question-answer pairs to construct a beneficial memory repository. Additionally, the FIFO Queue mechanism in MemGPT Packer et al. (2023) maintains a rolling history of messages, encapsulating interactions between agents and users, system notifications, and inputs and outputs of function calls.
@@
On the other hand, reflected rationales offer condensed explanations, such as rules, that support decision-making. For instance, TRAN Yang et al. (2023b) archives rules inferred from experiences alongside information on mistakes to mitigate future errors. Correspondingly, TiM Liu et al. (2023a) preserves inductive reasoning, defined as text elucidating the relationships between entities. Moreover, IML Wang et al. (2024a) and ICE Qian et al. (2024) store comprehensive notes and rules derived from a series of trajectories, demonstrating the broad spectrum of content types that memory systems can accommodate.
MemoryBank Zhong et al. (2024b) and AesopAgent Wang et al. (2024d) establish both experience and reflection knowledge stores, which are the integration of both two memories.
@@
Updating Operation: We categorize the operations to the memory into Insert, Reflect, and Forget. The most common operation is insert, methods insert text content into the memory for storage Li and Qiu (2023); Yang et al. (2023b); Zhong et al. (2024b); Packer et al. (2023); Liu et al. (2023a); Wang et al. (2024a). Another operation is reflection, which is to think and summarize previous experiences to conceptualize rules and knowledge for future use Yang et al. (2023b); Zhong et al. (2024b); Wang et al. (2024a); Qian et al. (2024). Last, due to the limited storage of memory, forgetting content is crucial to keeping memory efficient and the content valid. MemGPT Packer et al. (2023) adopts the FIFO queue to forget the contents. MemoryBank Zhong et al. (2024b) establishes a forgetting curve on the insert time of each item.
Working Memory
@@
The methods use past experience to evolve the capabilities of agents by updating internal memory streams, states, or beliefs, known as working memory, often in the form of verbal cues. Reflexion Shinn et al. (2023) introduces verbal reinforcement learning for decision-making improvement without conventional model updates. Similarly, IML Wang et al. (2024a) enables LLM-based agents to autonomously learn and adapt to their environment by summarizing, refining, and updating knowledge based on past experience directly in working memory.
@@
EvolutionaryAgent Li et al. (2024c) aligns agents with dynamically changing social norms through evolution and selection principles, leveraging environmental feedback for self-evolution. Agent-Pro Zhang et al. (2024d) employs policy-level reflection and optimization, allowing agents to adapt their behavior and beliefs in interactive scenarios based on past outcomes. Lastly, ProAgent Zhang et al. (2024a) enhances cooperation in multi-agent systems by dynamically interpreting teammates’ intentions and adapting behavior.
@@ 7 Evaluation
${\mathcal{E}}^{t+1},{\mathcal{S}}^{t+1}=f^{{\mathcal{E}}}(M^{t},{\mathcal{E}}^{t},\mathrm{ENV}),$ (7)
@@ 7.1 Quantitative Evaluation
This method focuses on providing measurable metrics to reliably assess LLM performance, such as automatic Papineni et al. (2002); Lin (2004) and human evaluation. However, traditional automatic metrics struggle to accurately evaluate increasingly complex tasks, and human assessment is not an ideal option for autonomous self-evolution. Recent trends use LLMs as human proxy for automatic evaluators, offering cost-effective and scalable solutions for evaluations.
For example, reward model score has been widely used to measure model or task performances Shinn et al. (2024) and select the best checkpoint Ouyang et al. (2022). LLM-as-a-judge Zheng et al. (2024a) using LLMs to evaluate LLMs, employing methods like pairwise comparison, single answer grading, and reference-guided grading. It shows that LLMs can closely match human judgment, enabling efficient large-scale evaluations.
@@ 7.2 Qualitative Evaluation
Qualitative evaluation involves case studies and analysis to derive insights, offering evolving guidance for subsequent iterations. Initiatives such as LLM-as-a-judge Zheng et al. (2024a) provide the reasoning behind its assessments; ChatEval Chan et al. (2023) explores the strengths and weaknesses of model outputs through debate mechanisms. Furthermore, TRAN Yang et al. (2023b) leverages past errors to formulate rules that enhance future LLM performances. Nonetheless, compared with instance-level critic or reflection, qualitative evaluation at the task- or model-level still needs comprehensive investigation.
@@ 8.1 Objectives: Diversity and Hierarchy
Section 3 summarizes existing evolution objectives and their coverage. Nonetheless, these highlighted objectives can only satisfy a small fraction of the vast human needs. The extensive application of LLMs across various tasks and industries highlights unresolved challenges in establishing self-evolution frameworks for evolving objectives that can comprehensively address a broader spectrum of real-world tasks Eloundou et al. (2023).
Furthermore, the concept of evolving objectives entails a potential hierarchical structure; for instance, UltraTool Huang et al. (2024b) and T-Eval Chen et al. (2023d) categorize the capability of tool usage into various sub-dimensions. Exploring evolutionary objectives into manageable sub-goals and pursuing them individually emerges as a viable strategy.
@@ 8.2 Level of Autonomy: From Low to High
As discussed in previous open problem (§ 8.1), there are a large of unfulfilled objectives. However, most of the existing self-evolution frameworks are at the Low-level which requires specifically designed modules Yuan et al. (2024); Lu et al. (2024a); Qiao et al. (2024). These frameworks are objective-dependent and rely on large human efforts to develop. Exhausting all objectives are not deployment-efficient which brings about the urgent need to develop medium and high levels self-evolution frameworks. At the medium level, it doesn’t require expert efforts to design specific modules. LLMs can self-evolve according to targeted objectives. Then at the high level, LLMs can investigate their current deficiencies and evolve in a targeted manner. In all, developing highly autonomous self-evolution frameworks remains an open problem.
@@ 8.3 Experience Acquisition and Refinement: From Empirical to Theoretical
Suppose we have addressed the previous two challenges that we have developed promising self-evolution frameworks, the exploration of self-evolution LLMs lacks solid theoretical grounding. This idea posits that LLMs can self-improve or correct their outputs, with or without feedback from the environment. However, the mechanisms behind it remain unclear. Studies show mixed results: Huang et al. (2023) observed self-corrective behavior in models with over 22 billion parameters, while Ganguli et al. (2023) finds LLMs struggle to self-correct reasoning errors without external feedback.
A related challenge is the use of self-generated data for learning. Critics argue this approach could reduce linguistic diversity Guo et al. (2023) and lead to "model collapse," where models fail to capture complex, long-tailed data distributions Shumailov et al. (2023). Furthermore, Alemohammad et al. (2023) reveal that generative models trained on their synthetic outputs progressively lose output quality and diversity. Fu et al. (2024) extend this by theoretically analyzing the impact of self-consuming training loops on model performance, emphasizing the importance of balancing synthetic and real data to mitigate error accumulation.
Recent studies Yang et al. (2024c); Singh et al. (2023) also show that current methods struggle to improve after more than three rounds of self-evolution. One hypothesized reason is that the self-critic of LLM has not co-evolved with the evolving objective, but more experimental and theoretical support is still needed. These findings highlight a pressing need for more theoretical exploration in self-evolving LLMs. Addressing these concerns is crucial for advancing the field and ensuring that models can effectively learn and improve over time.
@@ 8.5 Evaluation: Systematic and Evolving
To effectively assess LLMs, a dynamic, comprehensive benchmark is crucial. This becomes even more pivotal as we progress towards Artificial General Intelligence (AGI). Traditional static benchmarks risk obsolescence due to LLMs’ evolving nature and potential access to test data through interacting with environments, such as search engines, undermining their reliability. A dynamic benchmark, like Sotopia Zhou et al. (2023a), proposes a solution by creating an LLM-based environment tailored for evaluating the social intelligence of LLMs, thereby avoiding the limitations posed by static benchmarks.
@@ 8.6 Safety and Superalignment
The advancement of LLMs opens the possibility for AI systems to achieve or even surpass expert-level capabilities in both supportive and autonomous decision-making. For safety, ensuring these LLMs align with human values and preferences is crucial, particularly to mitigate inherent biases that can impact areas such as political debates, as highlighted by Taubenfeld et al. (2024). OpenAI’s initiative, Superalignment Leike and Sutskever (2023), aims to align a superintelligence by developing scalable training methods, validating models for alignment, and stress-testing the alignment process through scalable oversight Saunders et al. (2022), robustness Perez et al. (2022), automated interpretability Bills et al. (2023), and adversarial testing. Although challenges remain, Superalignment marks an initial attempt to develop a self-evolving LLM that closely aligns with human ethics and values in a scalable way.
@@ 9 Conclusion
Acknowledgements
This work was supported by Alibaba Group through Alibaba Research Intern Program.

### v2

@@ 6.2.1 External Memory
This approach utilizes an external module to collect, update, and retrieve past experiences and knowledge, enabling models to access a rich pool of insights and achieve better results without updating model parameters. The external memory mechanism is common in AI Agent systems [136, 89, 123]. This section provides a detailed overview of the latest methods for updating external memory, emphasizing the aspects of memory Content and Updating Operations, and summarized in Table 2.
@@
Content: External memory mainly stores two types of content: past experiences and reflected rationale, each serving distinct purposes. For instance, past experience provides valuable historical context, serving as a guiding force toward achieving improved outcomes. MoT [66] archives filtered question-answer pairs to construct a beneficial memory repository. Additionally, the FIFO Queue mechanism in MemGPT [84] maintains a rolling history of messages, encapsulating interactions between agents and users, system notifications, and inputs and outputs of function calls.
@@
On the other hand, reflected rationales offer condensed explanations, such as rules, that support decision-making. For instance, TRAN [140] archives rules inferred from experiences alongside information on mistakes to mitigate future errors. Correspondingly, TiM [72] preserves inductive reasoning, defined as text elucidating the relationships between entities. Moreover, IML [121] and ICE [89] store comprehensive notes and rules derived from a series of trajectories, demonstrating the broad spectrum of content types that memory systems can accommodate.
MemoryBank [161] and AesopAgent [123] reflect on experience and rationale, save them as external memory, and retrieve them when needed to achieve better performance.
@@
Updating Operation: We categorize the operations to the memory into Insert, Reflect, and Forget. The most common operation is insert, methods insert text content into the memory for storage [66, 140, 161, 84, 72, 121]. Another operation is reflection, which is to think and summarize previous experiences to conceptualize rules and knowledge for future use [140, 161, 121, 89]. Last, due to the limited storage of memory, forgetting content is crucial to keeping memory efficient and the content valid. MemGPT [84] adopts the FIFO queue to forget the contents. MemoryBank [161] establishes a forgetting curve on the insert time of each item.
6.2.2 Working Memory
@@
The methods use past experience to evolve the capabilities of agents by updating internal memory streams, states, or beliefs, known as working memory, often in the form of verbal cues. Reflexion [98] introduces verbal reinforcement learning for decision-making improvement without conventional model updates. Similarly, IML [121] enables LLM-based agents to autonomously learn and adapt to their environment by summarizing, refining, and updating knowledge based on past experience directly in working memory.
@@
EvolutionaryAgent [65] aligns agents with dynamically changing social norms through evolution and selection principles, leveraging environmental feedback for self-evolution. Agent-Pro [154] employs policy-level reflection and optimization, allowing agents to adapt their behavior and beliefs in interactive scenarios based on past outcomes. Lastly, ProAgent [151] enhances cooperation in multi-agent systems by dynamically interpreting teammates’ intentions and adapting behavior.
@@ 7 Evaluation
${\mathcal{E}}^{t+1},{\mathcal{S}}^{t}=f^{{\mathcal{E}}}(M^{t},{\mathcal{E}}^{t},\mathrm{ENV}),$ (7)
@@ 7.1 Quantitative Evaluation
This method focuses on providing measurable metrics to reliably assess LLM performance, such as automatic [86, 68] and human evaluation. However, traditional automatic metrics struggle to accurately evaluate increasingly complex tasks, and human assessment is not an ideal option for autonomous self-evolution. Recent trends use LLMs as human proxy for automatic evaluators, offering cost-effective and scalable solutions for evaluations.
For example, reward model score has been widely used to measure model or task performances [98] and select the best checkpoint [83]. LLM-as-a-judge [157] using LLMs to evaluate LLMs, employing methods like pairwise comparison, single answer grading, and reference-guided grading. It shows that LLMs can closely match human judgment, enabling efficient large-scale evaluations.
@@ 7.2 Qualitative Evaluation
Qualitative evaluation involves case studies and analysis to derive insights, offering evolving guidance for subsequent iterations. Initiatives such as LLM-as-a-judge [157] provide the reasoning behind its assessments; ChatEval [17] explores the strengths and weaknesses of model outputs through debate mechanisms. Furthermore, TRAN [140] leverages past errors to formulate rules that enhance future LLM performances. Nonetheless, compared with instance-level critic or reflection, qualitative evaluation at the task- or model-level still needs comprehensive investigation.
@@ 8.1 Objectives: Diversity and Hierarchy
Section 3 summarizes existing evolution objectives and their coverage. Nonetheless, these highlighted objectives can only satisfy a small fraction of the vast human needs. The extensive application of LLMs across various tasks and industries highlights unresolved challenges in establishing self-evolution frameworks for evolving objectives that can comprehensively address a broader spectrum of real-world tasks [35].
Furthermore, the concept of evolving objectives entails a potential hierarchical structure; for instance, UltraTool [54] and T-Eval [23] categorize the capability of tool usage into various sub-dimensions. Exploring evolutionary objectives into manageable sub-goals and pursuing them individually emerges as a viable strategy.
@@ 8.2 Level of Autonomy: From Low to High
As discussed in the previous open problem (§ 8.1), there are a large of unfulfilled objectives. However, most of the existing self-evolution frameworks are at the Low-level which requires specifically designed modules [148, 77, 90]. These frameworks are objective-dependent and rely on large human efforts to develop. Exhausting all objectives are not deployment-efficient which brings about the urgent need to develop medium and high levels self-evolution frameworks. At the medium level, it doesn’t require expert efforts to design specific modules. LLMs can self-evolve according to targeted objectives. Then at the high level, LLMs can investigate their current deficiencies and evolve in a targeted manner. In all, developing highly autonomous self-evolution frameworks remains an open problem.
@@ 8.3 Experience Acquisition and Refinement: From Empirical to Theoretical
Suppose we have addressed the previous two challenges and developed promising self-evolution frameworks, but the exploration of self-evolution LLMs still lacks solid theoretical grounding. This idea posits that LLMs can self-improve or correct their outputs, with or without feedback from the environment. However, the mechanisms behind it remain unclear. Studies show mixed results: Huang et al [51] observed self-corrective behavior in models with over 22 billion parameters, while Ganguli et al [38] finds LLMs struggle to self-correct reasoning errors without external feedback.
A related challenge is the use of self-generated data for learning. Critics argue this approach could reduce linguistic diversity [43] and lead to ”model collapse,” where models fail to capture complex, long-tailed data distributions [100]. Furthermore, Alemohammad et al [5] reveal that generative models trained on their synthetic outputs progressively lose output quality and diversity. Fu et al [37] extend this by theoretically analyzing the impact of self-consuming training loops on model performance, emphasizing the importance of balancing synthetic and real data to mitigate error accumulation.
Recent studies [141, 103] also show that current methods struggle to improve after more than three rounds of self-evolution. One hypothesized reason is that the self-critic of LLM has not co-evolved with the evolving objective, but more experimental and theoretical support is still needed. These findings highlight a pressing need for more theoretical exploration in self-evolving LLMs. Addressing these concerns is crucial for advancing the field and ensuring that models can effectively learn and improve over time.
@@ 8.5 Evaluation: Systematic and Evolving
To effectively assess LLMs, a dynamic, comprehensive benchmark is crucial. This becomes even more pivotal as we progress towards Artificial General Intelligence (AGI). Traditional static benchmarks risk obsolescence due to LLMs’ evolving nature and potential access to test data through interacting with environments, such as search engines, undermining their reliability. A dynamic benchmark, like Sotopia [163], proposes a solution by creating an LLM-based environment tailored for evaluating the social intelligence of LLMs, thereby avoiding the limitations posed by static benchmarks.
@@ 8.6 Safety and Superalignment
The advancement of LLMs opens the possibility for AI systems to achieve or even surpass expert-level capabilities in both supportive and autonomous decision-making. For safety, ensuring these LLMs align with human values and preferences is crucial, particularly to mitigate inherent biases that can impact areas such as political debates, as highlighted by Taubenfeld et al [112]. OpenAI’s initiative, Superalignment [61], aims to align a superintelligence by developing scalable training methods, validating models for alignment, and stress-testing the alignment process through scalable oversight [94], robustness [88], automated interpretability [11], and adversarial testing. Although challenges remain, Superalignment marks an initial attempt to develop a self-evolving LLM that closely aligns with human ethics and values in a scalable way.